EndersWorld:

The solution to the equation is:

EndersWorld:

\[4(x+2)=3(5-x) is:\] 7/15 1 13/7 7 15

dude:

Do you have an idea on how to solve this?

EndersWorld:

One second, gonna post my work, just need to make sure I did it right.

dude:

Oh okay

EndersWorld:

\[4(1+2)=3(5-1)\] 12=12

dude:

Looks right

EndersWorld:

I got a total of 17 questions, you got time?

dude:

Sure, make a new post =p

EndersWorld:

I’m also working through them as I post them, so may take a sec

dude:

I have time dw